OpenStudy (anonymous):

x^2 -3y^2 +8y =0 is that a hyperbola, ellipse or what? whats that in general form

OpenStudy (anonymous):

Hyperbola!

OpenStudy (lgbasallote):

hmm

OpenStudy (krabsen):

It is a hyperbola not a circle :)

OpenStudy (anonymous):

r is here radius...

OpenStudy (anonymous):

\[x^2 -3y^2 +8y=x^2 -3(y^2 -\frac{8}{3}y+\frac{16}{9}-\frac{16}{9})=x^2 -3(y^2 -\frac{8}{3}y+\frac{16}{9})+\frac{16}{3}=0\\ x^2 -3(y -\frac{4}{3})^2=-\frac{16}{3} \\ 3(y -\frac{4}{3})^2-x^2=\frac{16}{3}\] A hyperbola

OpenStudy (anonymous):

Equation of a hyperbola \[\frac{(y-h)^2}{a^2}-\frac{(x-k)^2}{b^2}=1\]